OpenStudy (anonymous):

what are all the real zeroes y=(x-12)^3-7

OpenStudy (anonymous):

help @Hero

OpenStudy (anonymous):

please!

OpenStudy (anonymous):

Hello???????

hero (hero):

Hint: set y = 0 and solve for x

OpenStudy (anonymous):

so 0=(x-12)^3-7 0=x^3-12^3-7 0=x^3-1728-7 0=x^3-1721

OpenStudy (anonymous):

is this right

OpenStudy (anonymous):

@Hero is this going on the right track

OpenStudy (anonymous):

HELLLOOOO???? ANYONE!!! I NEED HELP!

OpenStudy (anonymous):

@yummydum can u help

OpenStudy (anonymous):

Hello? @yummydum

OpenStudy (tamtoan):

i don't think hero do the expansion of (x-12)^3 right.....should move the number to the other side and take cube root of both side

OpenStudy (tamtoan):

yummydum, why do you use the method of finding roots for quaratic equation to solve a third degree equation?

OpenStudy (anonymous):

can u help @tamtoan

OpenStudy (anonymous):

i mean this is a pretty icky equation they've got here

OpenStudy (tamtoan):

@lala2, i just typed earlier , as hero said, set y = 0 then you have (x-12)^3 - 7 = 0. ==> (x-12)^3 = 7 (x-12) =  then continue from here

OpenStudy (anonymous):

whats after that

OpenStudy (tamtoan):

@yummydum , you should never use the formula of solving quadratic equation and apply it to the third degree equation unless in some special case you can reduce to second degoree equation, otherwise, the answer will never be right

OpenStudy (anonymous):

im confused

OpenStudy (tamtoan):

 and you can't solve this ?